# Toward the Categorical Data Map

### v3

@@ Abstract
Categorical data does not have an intrinsic definition of distance or order, and therefore, established visualization techniques for categorical data only allow for a set-based or frequency-based analysis, e.g., through Euler diagrams or Parallel Sets, and do not support a similarity-based analysis. We present a novel dimensionality reduction-based visualization for categorical data, which is based on defining the distance of two data items as the number of varying attributes. Our technique enables users to pre-attentively detect groups of similar data items and observe the properties of the projection, such as attributes strongly influencing the embedding. Our prototype visually encodes data properties in an enhanced scatterplot-like visualization, encoding attributes in the background to show the distribution of categories. In addition, we propose two graph-based measures to quantify the plot’s visual quality, which rank attributes according to their contribution to cluster cohesion. To demonstrate the capabilities of our similarity-based approach, we compare it to Euler diagrams and Parallel Sets regarding visual scalability and show its benefits through an expert study with five data scientists analyzing the Titanic and Mushroom datasets with up to 23 attributes and 8124 category combinations. Our results indicate that the Categorical Data Map offers an effective analysis method, especially for large datasets with a high number of category combinations.
Index Terms:
@@
I Introduction
@@
Categorical data can be encountered in numerous domains, such as representing inventory data describing product properties like color in sales or bioinformatics, encoding the genes formed by nucleotide sequences. In contrast to numeric and ordinal data, categorical data does not have an intrinsic order or distance associated with each value pair. The visual analysis of categorical data is challenging since categorical data describes an attribute by name only, with the only supported operators being equality and mode.
@@
Currently, there are two widespread methods of visualizing categorical data: (1) Frequency-based visualizations [1, 2, 3] map the categorical values to their frequencies, for example, through bar charts, pie charts, or enhanced variants, such as stacked bar charts. In contrast, (2) set visualizations solely focus on the set nature of categorical data items, specifically their intersections [4], such as Euler diagrams (see Fig. 1 (left)). Set visualizations like Euler diagrams do not scale well for sets with many intersections because visual clutter is detrimental to their readability. Other, less common solutions treat dimensions independently and map data to a continuous design model [5, 6, 7], leveraging visualization types that initially have been designed for numerical data, such as scatterplots or parallel coordinate plots. However, these approaches deviate from the discrete nature of categorical data and suffer from visual clutter and overplotting, limiting their readability [8]. Approaches, such as Parallel Sets [9] and Sankey diagrams [10], follow the frequency and set-based paradigms (see Fig. 1 (right)). These approaches trade effectiveness in visualizing the presence of subsets for the presentation of frequency information. Thus, these approaches require additional design considerations since they tend to emphasize preselected attributes over others [11].
Figure 1: Examples of existing categorical data visualizations: An Euler diagram of Harry Potter characters [12] (left) and overlap reduced Parallel Sets [11, 9] visualization of the Property Sales dataset [13] (right).
@@
None of the previously described techniques support the distance-based analysis of categorical data, i.e., encoding the similarity of categorical data items using distance such that similar data items are placed close to each other while differing data items are positioned far apart. Analyzing categorical data based on a group or subset similarity is useful, e.g., visually clustering data items only differing in a few attributes can help us better understand important characteristics of the group. Generally, this would allow us to apply methods from cluster analysis to categorical data.
@@
Thus, the challenge is finding an effective way to describe distance relations between categorical data items to represent the dissimilarity of data points such that the distance reflects their dissimilarity. At the same time, the whole visualization remains coherent and interpretable. We further need to find a visual mapping that enables the interpretation of distances and simultaneously conveys the properties of data items, i.e., effectively visualizing an item’s attributes by using color and position to visually encode attributes. Through this, We address the combinatorial problem of categorical data, i.e., that with the increasing number of attributes and categories, the number of required colors to represent a category with distinguishable colors becomes increasingly difficult.
Tacking these challenges, we contribute the following:
@@
    A design study using 2D-projection methods, i.e., dimensionality reduction techniques, for categorical data while visually encoding the category distribution into the background. Through layout enrichment, we enable the exploration of the data distribution, enhancing orientation and navigation. Additionally, we contribute four glyph designs to represent categorical subsets.
  2. (2)
    A description and discussion of distance measures for categorical data based on the interpretation of data items as sets and one-hot encoding, which allows representing dissimilarity as the number of differing elements.
@@
    An online demonstrator (dennig.dbvis.de/categorical-data-map) making the acquired results accessible. To further aid reproducibility, we openly publish all our datasets and source code via OSF (osf.io/jzd46) and the Data Repository of the University of Stuttgart (DaRUS)  [14].
@@
II Related Work
II-A Visualization Techniques for Categorical Data
@@
Set visualization is one of the core techniques to for categorical data. To visualize the members of sets and their intersections, Venn and Euler diagrams are the two most prevalent representations [15]. Multiple adaptations of both techniques mitigate challenges, e.g., to preserve semantics [16], draw area-proportional diagrams [17], or incorporate glyphs to show additional information [18]. Other set visualization techniques use lines to indicate set intersections [19] and matrices to show the cardinality of intersection sets [20], or include the semantic context to visualize sets [21]. Alsallakh et al. presented a comprehensive survey on set visualizations [4]. There are also frequency-based visualization methods that focus on attribute frequencies, such as Mosaic plots [1] and Parallel Bargrams [2] by mapping data item occurrences to one or multiple attributes, e.g., a rectangle’s area. Other methods map data to a continuous design model, such that they are compatible with visualization for numeric data, e.g., Rosario et al. [7] describe the mapping of categorical data to numeric values for the visualization in Parallel Coordinates [22]. Hybrid methods consider both aspects, e.g., Parallel Sets [9, 8] and Sankey diagrams [10]. However, Parallel Sets and Sankey diagrams can suffer from the Müller-Lyer and Sine illusions [23, 24] where lines seem to vary in distance or length.
@@
While plenty of approaches visualize categorical data, to the best of our knowledge, none focus on similarity analysis.
II-B Dimensionality Reduction for Categorical Data
@@
Our approach makes use of dimensionality reduction. However, there exist dimensionality reduction methods for categorical data that do not focus on similarity but rather describe the central oppositions in the data [25]. When needing to reduce the dimensionality of categorical data, Correspondence Analysis (CA), similar to Principal Component Analysis (PCA) [26] for numerical data, extracts the standard coordinates, yielding a Biplot [27] of the reduced space. In case of more than two categorical variables, Multiple Correspondence Analysis (MCA) can be used to reduce the number of dimensions showing the central oppositions [25]. Factor analysis of mixed data (FAMD) is a principal component technique for continuous and categorical variables [28]. The continuous variables are scaled to unit variance, and the categorical variables are transformed into a disjunctive data table and then scaled using the specific scaling of MCA to balance the influence of both continuous and categorical variables in the analysis. Multiple Factor Analysis (MFA) combines these methods for mixed data: It uses PCA when variables are quantitative, MCA when variables are qualitative, and FAMD when the active variables belong to either of the two types. The Data Context Map [29] visualizes mixed-data using an MDS-based plot and displaying categorical attributes on top of the projection while also coloring points and regions according to the predominant category. The approach by Thane et al. [30] uses force-directed graph layouts to visualize categorical datasets representing categories as nodes while edges represent their co-occurrence.
In contrast to existing methods, we derive a similarity relation for categorical data by first aggregating data points and subsequently defining a distance between the subsets.
II-C Layout Enrichment for 2-Dimensional Data Projections
@@
The idea to enrich scatterplot layouts by encoding additional information in the background of a projection is not new [31]. The main usage occurs for the visualization of distortions in the topology of the embedding resulting from dimensionality reduction [32]. Lespintas and Aupetit proposed CheckViz [33], visualizing the presence of tears (i.e., missing neighborhood) and shuffled data (i.e., wrong neighborhood). Broeksema et al. explored the visualization of categorical data, combining MCA with an enhanced treeview to integrate data record information. They used dimensionality reduction to plot categorical data and showed. However, they did not address the high redundancy of categorical datasets [34]. Sohns et al. followed a similar approach; however, they used non-linear dimensionality reduction methods to project mixed data while using categorical attributes to highlight areas of the embedding space. However, this approach excludes all categorical attributes from the dimensionality reduction process altogether [35]. Morariu et al. encode the projection’s quality into the background of scatterplots [36] while projecting projections.
In this work, we use layout enrichment to visualize the layout of different categories of attributes in dimensionality reduction, allowing users to observe the distribution and grouping of data items and subsets using Voronoi diagrams.
II-D Measures for Quality and Patterns in Visualizations
@@
Quality measures for visualizations describe a set of measurements designed to optimize visualizations in terms of readability and clutter reduction [37]. Other measures quantify the presence of patterns in a visualization. Instead of removing any information from the visualization, metrics measure their characteristics, which can be used to compare and rank different visualizations. Examples are: Magnostics for matrix visualizations [38], Scagnostics for general patterns and trends on scatterplots of numeric data [39], SepMe, a machine-learning based approach to quantify the presence of clusters in scatterplots [40], ClustMe quantifies the visual separation of classes in scatterplots [41], Pargnostics for parallel coordinate plots [42], Visualgnostics projections of high-dimensional data [43], Pixgnostics for pixel-based visualizations [44], and ParSetgnostics for Parallel Sets [11].
@@
We contribute two graph-based measures for the quantification of visual quality for 2-dimensional projections of categorical data. In this way, we improve the exploration of categorical data by recommending layout-enriched views according to their visual structure.
Figure 2: The Categorical Data Map enables projection-based analysis of categorical data. The analysis of Property Sales [13] with t-SNE [45] using the Overlap coefficient [46]; (1) shows eight clusters, created by the attributes Purchaser Currently Living In (split left-right) and Location of Purchased Property (four symmetric clusters on each half). To show the category distribution, we visualize the categories of attributes in (2)-(4). (2) shows a clear boundary between Purchasers Living In Private Property or Public Property. (3) Clear boundaries are also visible for the Location of Property attribute, while in (4) the Property Type Purchased contributes least to the clusters. Considering the glyph sizes, the categories Private Propriety and Central occur often together.
III Constructing the Categorical Data Map
@@
Typically, categorical datasets exhibit inherent sparsity, i.e., only a fraction of all possible combinations category combinations is present in a dataset, e.g., for the Mushroom dataset, only 8124 out of 243.799.621.632.000 possible combinations. Thus, we assume that there are relationships among the existing categories restricting their distribution. Additionally, categorical datasets can be highly redundant, e.g., the Titanic dataset contains 2201 data entries but only 24 unique entries, i.e., only 24 subsets, considering all attributes. Thus, we leverage these properties in the design of the Categorical Data Map as an analytical approach for the similarity-based analysis of categorical subsets with the following requirements:
  1. (R1)
@@
    Distance of categorical subset in a scatterplot should indicate similarity, i.e., subsets with a low distance should differ in fewer attributes than subsets with a high distance.
  2. (R2)
@@
An example of our approach is shown in Fig. 2. (R1) and (R2) are described further in subsection III-A and subsection III-B. We address (R3) by evaluating different glyph designs for subsets of categorical data (see subsection III-C). We address (R4) in subsection III-E, describing measures to rank attributes according to their degree of splitting the embedding into connected areas. In the following, we describe how we derive distance relations of categorical data and how a projection-based approach, i.e., the Categorical Data Map, is constructed.
III-A Data Representations and Distance Measures
In this section, we describe data representations and distance measures that we use to project categorical subsets using dimensionality reduction. In the following, we define the relationship between data items and subsets to address (R1) and (R2). We selected set-based and numeric representations and distance measures based on their common use in visual data analysis [47] and highlight relations between them and other representations where applicable. In the following, we describe data abstractions for categorical data.
Set Representation: A data item is represented by a set with the same size as the number of attributes of the dataset. In the set representation, every category for every attribute gets a unique symbol, representing its value. The relation between the representative sets is used to determine the relation between two data items and enables set operations. We use the union and the intersection set operation to define distance relations.
Regular One-Hot Encoding: The conversion of the attributes of a data item into a numeric representation can be achieved via one-hot encoding, often used in machine learning [48, 49]. The result is a binary vector where the presence of an attribute is displayed by an on-bit at a specific position. Subsequently, distance measures for numeric data, such as Manhattan or Euclidean distance, can be applied.
Reduced One-Hot Encoding: In addition, we construct a reduced version of one-hot encoding, where binary attributes, i.e., attributes with only two possible categories, are not transferred into two dimensions but a single one. This leads to non-binary categories being weighted higher than binary ones in the projection. This kind of distortion needs to be considered when using this representation.
We use the representation described above, with the corresponding type of distance measure defined in the following. Distance measures for sets are used with set-based representations, while distance measures for numeric representations are used with one-hot encoding-based representations.
Distance Measures for Sets: Our approach uses existing distance measures that can be adapted to categorical data. We choose these measures based on the recommendations and examples from Cha’s comprehensive survey on distance functions [50], as well as the evaluations of similarity measures for categorical data [51, 52], intending to model similarity for vector representations and set representations. We found three set-based similarity coefficients describing the similarity of sets. In the following, we define $A$ and $B$ as two sets.
One measure for the similarity of sets is the Overlap coefficient, which is defined as
$O(A,B)\coloneqq\left|A\cap B\right|/\min\{\left|A\right|,\left|B\right|\}$ (1)
and its distance $d_{O}(A,B)\coloneqq 1-{O}(A,B)$ [46]. Akin to the Hemming distance [53], which measures how many positions differ when comparing two strings, the Overlap coefficient gives us the ratio of the size of the intersection relative to the size of the smaller of the two. Measuring the number of differing attributes is easier to understand then the following measures, since it does not add additional emphasis.
The Jaccard Similarity Index is defined as
$JSI(A,B)\coloneqq\left|A\cap B\right|/\left|A\cup B\right|$ (2)
and its associated distance or dissimilarity measure $d_{JSI}(A,B)\coloneqq 1-{JSI}(A,B)$ [54]. Jaccard Similarity Index differs from the Overlap coefficient in that it measures the number of elements in the intersection of the two sets relative to their union, which, in comparison to the Jaccard distance, will yield a lower similarity and a larger distance for differing sets, and thus, provide bigger effect on dissimilar sets. This effect needs to be considered when using this distance measure.
The Sørensen-Dice coefficient is defined as
$SD(A,B)\coloneqq(2\,\left|A\cap B\right|)/(|A|+|B|)$ (3)
and its distance $d_{SD}(A,B)\coloneqq 1-SD(A,B)$ [55, 56]. It defines a semi-metric, since it does not full fill the requirements of the triangle inequality.
Distance Measures for Numeric Representations: For determining distances between one-hot encoded representations, we use two Minkowski distance measures. In the following, we define $X$ and $Y$ as two vectors of equal size. The Minkowski distance is defined as
$d_{p}(X,Y)\coloneqq\sqrt[p]{\sum\limits_{i=1}^{|D|}\left|X_{i}-Y_{i}\right|^{p}}\text{ with }p\in\mathbb{N}$ (4)
The survey on distance functions by Cha [50] describes the Minkowski distance functions in general but only gives practical application for the Euclidean and the Manhattan distance. Therefore, we limit ourselves to those two. The Euclidean distance, which is defined as $d_{p}$ with $p=2$. The Euclidean distance enforces a spacial and geometric relationship, i.e., as its name suggests, the distance relationship points have in N-dimensional Euclidean space for N-dimensional input vectors. The Manhattan distance is defined as $d_{p}$ with $p=1$. The Manhattan distance treats the dimensions of the input vectors as independent or orthogonal to each other.
In literature, there are also distance measures for mixed data, such as the Gower distance [57] or the “unified metric for the clustering of categorical and numerical attributes” by Cheung and Jia [58]. These measures combine distances for categorical, categorical, and numeric data, e.g., the Gower distance uses the Manhattan distance for numeric and ordinal attributes and the Sørensen-Dice coefficient for categorical attributes. We exclude these measures since they work with mixed data, which, given the preprocessing described in the previous subsection, will automatically act like the Sørensen-Dice for set-based representations or the Manhatten distance for on-hot encoding-based representations.
Relationships Between Distance Measures: For our distance calculations, we can assume that the sets representing a categorical data entire all have the same size, i.e., $|A|=|B|$. This has an impact on the result of distance measures, leading to identical results. With this assumption, the Overlap and Sørensen-Dice coefficients act identical, i.e., $SD(A,B)\equiv O(A,B)$ for $|A|=|B|$ (see Proof 1 in the supplementary material).
Additionally, the Manhattan distance acts like the Overlap coefficient if used with classical one-hot encoding (OHE). In OHE, the absence and presence of a category are represented as one bit of information. As the input to the Manhattan distance, this yields the number of bits that differ between the two inputs. This is identical to the simplified dissimilarity measure (i.e., distance measure) of the Overlap coefficient, which can be simplified to $d_{O}^{simp}(A,B)=|A|-|A\cap B|$, defining a metric (see Proof 2 in the supplementary material).
However, for the reduced form of OHE, the quality of Manhattan distance to any other distance or similarity no longer holds. This transformation gives additional weight to binary attributes by representing two categories with only one bit, skipping the representation of the absence or presence of the complementary category. The Jaccard Similarity Index and the Euclidean distance have unique behavior in all cases.
Thus, we do not need to consider all possible combination when using different data representations and distance measures with the Categorical Data Map.
III-B Projecting Categorical Data

### v4

# The Categorical Data Map: A Multidimensional Scaling-Based Approach

@@ Abstract
Categorical data does not have an intrinsic definition of distance or order, and thus, established visualization techniques for categorical data only allow for a set-based or frequency-based analysis, e.g., through Euler diagrams or Parallel Sets, and do not support a similarity-based analysis. We present a dimensionality reduction-based visualization for categorical data based on defining the distance of two data items as the number of varying attributes. Our technique enables users to pre-attentively detect groups of similar data items and observe the properties of the projection, such as attributes strongly influencing the embedding. Our prototype visually encodes data properties in an enhanced scatterplot-like visualization, visualizing attributes in the background to show the distribution of categories. We propose two graph-based measures to quantify the plot’s visual quality for ranking attributes according to their contribution to cluster cohesion. To demonstrate the capabilities of our method, we compare it to Euler diagrams and Parallel Sets regarding visual scalability and evaluate it quantitatively on seven real-world datasets using a range of common quality metrics. We conducted an expert study with five data scientists analyzing the Titanic and Mushroom datasets with up to 23 attributes and 8124 category combinations. Our results indicate that our Categorical Data Map is an effective analysis method for large datasets with a high number of category combinations.
keywords:
@@
[…]
The Categorical Data Map enables projection-based analysis of categorical data here exemplified by the Property Sales dataset [48] with MDS [80] using the Jaccard coefficient [39]: (1) shows 10 groups without layout enrichment. Our method reveals the patterns annotated in (1) in plots (2)-(4). (2) shows a clear separation between Private Property vs Public Property. (3) indicates boundaries and symmetries for the Location of Purchased Property attribute, while in (4), the Property Type Purchased contributes the least to the clusters. The glyph sizes encode the subset sizes, revealing that categories Private Propriety and Central often occur together.
Introduction
@@
Categorical data can be encountered in numerous domains, such as representing inventory data describing product properties like color in sales or bioinformatics, encoding the genes formed by nucleotide sequences [2]. In contrast to numeric and ordinal data, categorical data does not have an intrinsic order or distance associated with each value pair. The visual analysis of categorical data is challenging since categorical data describes an attribute by name only, with the only supported operators being equality, set membership, and mode.
@@
Currently, there are two widespread methods of visualizing categorical data: (1) Frequency-based visualizations [37, 85, 75] map the categorical values to their frequencies, for example, through bar charts, pie charts, or enhanced variants, such as stacked bar charts. In contrast, (2) set visualizations solely focus on the set nature of categorical data items, specifically their intersections [4]. Examples include such as Euler diagrams [62] and UpSet plots [53]. Set visualizations like Euler diagrams do not scale well for sets with many intersections because visual clutter is detrimental to their readability. Other, less common solutions treat dimensions independently and map data to a continuous design model [78, 40, 68], leveraging visualization types that initially have been designed for numerical data, such as scatterplots or parallel coordinate plots. However, these approaches deviate from the discrete nature of categorical data and suffer from visual clutter and overplotting, limiting their readability [49]. Approaches, such as Parallel Sets [15] and Sankey diagrams [46], follow the frequency and set-based paradigms. These approaches trade effectiveness in visualizing the presence of small subsets for the presentation of frequency information. These approaches require additional design considerations since they tend to emphasize preselected attributes over others [28].
@@
None of the previously described techniques support the similarity-based analysis of categorical data, i.e., deriving the similarity of categorical data items as distances such that similar data items are placed close to each other while differing data items are positioned far apart. Analyzing categorical data based on a group or subset similarity is useful, e.g., visually clustering data items only differing in a few attributes can help us better understand important characteristics of the group. Generally, this would allow us to apply methods from cluster analysis to categorical data.
@@
We follow the suggestion by Broeksema et al. [19] to investigate multidimensional scaling to generate visual mappings that enable the interpretation of distances and simultaneously convey the properties of data items, i.e., effectively visualizing an item’s attributes by using color and position to visually encode attributes. Through this, we address the combinatorial problem of categorical data, i.e., that with the increasing number of attributes and categories, the number of required colors to represent a category with distinguishable colors becomes increasingly difficult. Tackling these challenges, we contribute the following:
@@
    A technique applying multidimensional scaling to categorical data while visually encoding the category distribution into the background. Through layout enrichment, we enable the exploration of the category distribution, enhancing orientation and navigation. Additionally, we contribute four glyph designs to represent categorical subsets.
@@
    A quantitative comparison to multiple correspondence analysis-based projections and a qualitative expert study validating the effectiveness of our approach.
  4. (4)
    An online demonstrator (https://dennig.dbvis.de/categorical-data-map) making the acquired results accessible. To further aid reproducibility, we openly publish all our datasets and source code via OSF (osf.io/jzd46) and the Data Repository of the University of Stuttgart (DaRUS) [29].
@@
1 Related Work
Our approach is related to visualization and dimensionality reduction methods for categorical data. Furthermore, we propose a layout enrichment for multidimensional projections and contribute visual quality metrics for categorical data projections.
1.1 Visualization Techniques for Categorical Data
@@
Set visualization is one of the core techniques for categorical data. To visualize the members of sets and their intersections, Venn and Euler diagrams are the two most prevalent representations [12]. Multiple adaptations of both techniques mitigate challenges, e.g., to preserve semantics [44], draw area-proportional diagrams [65], or incorporate glyphs to show additional information [58]. Other set visualization techniques use lines to indicate set intersections [66] and matrices to show the cardinality of intersection sets [54], or include the semantic context to visualize sets [57]. Alsallakh et al. presented a comprehensive survey on set visualizations [4]. There are also frequency-based visualization methods that focus on attribute frequencies, such as Mosaic plots [37] and Parallel Bargrams [85] by mapping data item occurrences to one or multiple attributes, e.g., a rectangle’s area. Other methods map data to a continuous design model, such that they are compatible with visualization for numeric data, e.g., Rosario et al. [68] describe the mapping of categorical data to numeric values for the visualization in Parallel Coordinates [38]. Hybrid methods consider both aspects, e.g., Parallel Sets [15, 49] and Sankey diagrams [46]. However, Parallel Sets and Sankey diagrams can suffer from the Müller-Lyer and Sine illusions [27, 81] where lines seem to vary in distance or length, affecting the accurate interpretation of frequencies and proportions.
@@
While plenty of approaches visualize categorical data, to the best of our knowledge, none allows identifying groups of similar data items. Thus, we propose a visualization that focuses on similarity.
1.2 Dimensionality Reduction for Categorical Data
@@
Our approach makes use of dimensionality reduction (DR). However, there exist DR methods for categorical data that do not focus on similarity but rather describe the central oppositions in the data [34]. When needing to reduce the dimensionality of categorical data, Correspondence Analysis (CA), similar to Principal Component Analysis (PCA) [42] for numerical data, extracts the standard coordinates, yielding a Biplot [33] of the reduced space. In case of more than two categorical variables, Multiple Correspondence Analysis (MCA) can be used to reduce the number of dimensions showing the central oppositions [34]. Factor analysis of mixed data (FAMD) is a principal component technique for continuous and categorical variables [63]. The continuous variables are scaled to unit variance, and the categorical variables are transformed into a disjunctive data table and then scaled using the specific scaling of MCA to balance the influence of both continuous and categorical variables in the analysis. Multiple Factor Analysis (MFA) combines these methods for mixed data: It uses PCA when variables are quantitative, MCA when variables are qualitative, and FAMD when the active variables belong to both of the two types. The Data Context Map [23] visualizes mixed-data using an MDS-based plot and displays categorical attributes on top of the projection while also coloring points and regions according to the predominant category. The approach by Thane et al. [79] uses force-directed graph layouts to visualize categorical datasets representing categories as nodes while edges represent their co-occurrence.
MCA can embed categorical data but, like PCA, is a linear dimensionality reduction technique and thus not able to detect non-linear relationships [34, 19]. We propose using MDS to visualize the similarity of categorical data points in a scatterplot-like layout.
1.3 Layout Enrichment for 2-Dimensional Data Projections
@@
The idea to enrich scatterplot layouts by encoding additional information in the background of a projection is not new [61]. The main usage occurs for the visualization of distortions in the topology of the embedding resulting from DR [6]. The following approaches make use of Voronoi diagrams [9] to encode additional information in the background of a projection. Lespinats and Aupetit proposed CheckViz [52], visualizing the presence of tears (i.e., missing neighborhood) and shuffled data (i.e., wrong neighborhood). Broeksema et al. explored the visualization of categorical data, combining MCA with an enhanced treeview to integrate data record information visualizing user-selected categories. However, they did not address the high redundancy of categorical datasets [19]. Sohns et al. followed a similar approach; however, they used non-linear DR methods to project mixed data while using categorical attributes to highlight areas of the embedding space. However, this approach excludes all categorical attributes from the DR process altogether [72]. DICON enables the analysis of multidimensional clusters with an interactive icon-based visualization that encodes additional statistical information visually using space-filling methods, including Voronoi diagrams [21]. Aside from using Voronoi diagrams, other methods for layout enrichment exist [16]. Morariu et al. encode the projection’s quality into the plot’s background using contours showing the embedding of projections called the metamap [60].
Layout enrichment methods largely focus on visualizing distortions of the projection. The approach by Broeksema et al. [19] does not address the analysis of a single attribute, so we propose a new enrichment that encodes the category of an attribute using color.
1.4 Metrics for Quality and Patterns in Visualizations
@@
Quality metrics for visualizations describe a set of measurements designed to optimize visualizations in terms of readability and clutter reduction [14]. Other metrics quantify the presence of patterns in a visualization. Instead of measuring quality, pattern metrics can be used to compare and rank different visualizations based on specific properties. Examples are: Magnostics for matrix visualizations [13], Scagnostics for general patterns and trends on scatterplots of numeric data [84], Pargnostics for parallel coordinate plots [25], Visualgnostics for projections of high-dimensional data [51], Pixgnostics for pixel-based visualizations [70], and ParSetgnostics for Parallel Sets [28]. SepMe is a machine-learning-based approach to quantify the presence of clusters in scatterplots [8], while ClustMe quantifies the visual separation of classes in scatterplots [1]. Aupetit and Catz [7] addressed the analysis of high-dimensional labeled data using graphs, including Voronoi diagrams. However, this approach does not address categorical data analysis, i.e., where no numerical attributes are present.
@@
We contribute two novel measures for quantifying visual quality for 2-dimensional projections of categorical data. In this way, we improve the exploration of categorical data by recommending layout-enriched views according to their visual structure.
2 Constructing the Categorical Data Map
@@
Typically, categorical datasets exhibit inherent sparsity, i.e., only a fraction of all possible category combinations is present in a dataset, e.g., for the Mushroom dataset, only 8124 out of 243.799.621.632.000 possible combinations. Thus, we assume that there are relationships among the existing categories restricting their combinations. Additionally, categorical datasets can be highly redundant, e.g., the Titanic dataset contains 2201 data items but only 24 unique entries, i.e., all data items can be assigned to one of 24 subsets. Thus, we focus on categorical subsets as subsets of unique attribute values. These subsets are our main representations, enabling us to assign a frequency. We leverage these properties in the design of the Categorical Data Map as an analytical approach for the similarity-based analysis of categorical subsets with the following constraints:
  1. (C1)
@@
    Distances of categorical subsets in a scatterplot should indicate similarity, i.e., subsets with a smaller distance should differ in fewer attributes than subsets with a larger distance.
  2. (C2)
@@
An example of our approach is shown in The Categorical Data Map: A Multidimensional Scaling-Based Approach. (C1) and (C2) are described further in Sec. 2.1. We address (C3) by evaluating different glyph designs and layout enrichments for subsets of categorical data (see Sec. 2.2). We address (C4) in Sec. 2.3, describing measures to rank attributes according to their degree of splitting the embedding into connected areas. In the following, we describe how we derive distance relations of categorical data and how a projection-based approach, i.e., the Categorical Data Map, is constructed.
2.1 Projecting Categorical Data